OpenStudy (anonymous):

How do you convert y-3=-2(x+1)^2?

OpenStudy (anonymous):

you meant simplify ?

OpenStudy (gorv):

y=3-2(x^2+1+2x) y=3-2x^2-2-4x y=-2x^2-4x+1

OpenStudy (anonymous):

I'm sorry i meant how do you convert y-3=-2(x+1)^2 into standard form

OpenStudy (ranga):

Just expand the terms by multiplying. Expand (x + 1)^2 first. Multiply it by -2 and then add 3 to it and simplify. The standard form is: y = ax^2 + bx + c

OpenStudy (anonymous):

thank you :)

OpenStudy (ranga):

you are welcome. And what is the answer you are getting?

OpenStudy (anonymous):

y=-2x^2-4x+1